### v2

# Formalism 25

Abstract.
Abraham Robinson’s philosophical stance has been the subject of several recent studies. Erhardt following Gaifman claims that Robinson was a finitist, and that there is a tension between his philosophical position and his actual mathematical output. We present evidence in Robinson’s writing that he is more accurately described as adhering to the philosophical approach of Formalism. Furthermore, we show that Robinson explicitly argued against certain finitist positions in his philosophical writings. There is no tension between Robinson’s mathematical work and his philosophy because mathematics and metamathematics are distinct fields: Robinson advocates finitism for metamathematics but no such restriction for mathematics. We show that Erhardt’s analysis is marred by historical errors, by routine conflation of the generic and the technical meaning of several key terms, and by a philosophical parti pris. Robinson’s Formalism remains a viable alternative to mathematical Platonism.
Key words and phrases:
Finitism; Formalism; foundations; Platonism; Realism; Robinson
2020 Mathematics Subject Classification:
Primary 00A30 Secondary 01A60, 03A05
March 16, 2025
1. Formalism and finitism
Abraham Robinson’s intellectual profile continues to inspire passions fifty years after his passing. A number of publications over the past few decades have analyzed his philosophical position, including [Benis-Sinaceur 1988, p. 603], [Gaifman 2012], [Sherry and Katz 2012], [Sanders 2020], [Weir 2024], and most recently [Erhardt 2024]. A majority of commentators take it for granted that Robinson’s mature position is best described as mathematical Formalism. However, Erhardt reads Robinson as a finitist (see Section 1.2 for a discussion of distinct meanings of the term) who rejects infinite totalities.
1.1. Meaning and reference
Significantly, there is a key observation in Robinson’s 1975 article that is missing from Erhardt’s text (though he does cite the article). The observation helps understand Robinson’s take on meaning. In 1975, Robinson clarified his position by stating that
mathematical theories that, allegedly, deal with infinite totalities have no detailed meaning, i.e. reference. [Robinson 1975, p. 42; emphasis added]; reprinted in [Robinson 1979, p. 557]
Namely, the term infinite totality has no reference (or referent) in either the physical world or any Platonic realm of mathematical abstracta. Robinson’s main goal here was to distance himself from mathematical Platonism. He did not believe that expressions such as infinite totality lacked meaning in the sense of being ‘pointless’ or ‘devoid of significance’, as we will see below; he only intended that such expressions lacked a reference, as we explained. In his earlier philosophical text Formalism 64 he used the term direct interpretation in place of reference:
I … regard a theory which refers to an infinite totality as meaningless in the sense that its terms and sentences cannot possess the direct interpretation in an actual structure that we should expect them to have by analogy with concrete (e.g., empirical) situations.111[Robinson 1965, p. 231]; emphasis on ‘meaningless’ in the original; emphasis on ‘direct interpretation’ added.
Having outlined Robinson’s position, let us examine Erhardt’s presentation thereof. Erhardt misrepresents Robinson’s position by quoting Robinson out of context. Thus, Erhardt claims the following:
In Formalism 64, Robinson’s most fully developed philosophical paper, he writes (Robinson 1964/1979b, pp. 230–231, italics in the original), “[…] the notion of a particular class of five elements, e.g. of five particular chairs, presents itself to my mind as clearly as the notion of a single individual […]. By contrast, I feel quite unable to grasp the idea of an actual infinite totality. To me there appears to exist an unbridgeable gulf between sets or structures of one, or two, or five elements, on one hand, and infinite structures on the other hand […]” [^2] Robinson as quoted in [Erhardt 2024, p. 2]; emphasis on “unable to grasp” added.
We stress the profusion of ellipses ‘[…]’ in Erhardt’s quotation from Robinson. Superficially, the abridged passage may suggest a finitist position. It is instructive to explore the issue whether Erhardt’s deletions alter the meaning of the passage as intended by Robinson. On the face of it, the passage as quoted by Erhardt sounds rather odd. A modern mathematician who claims to be “unable to grasp” the concept of an actual infinite totality certainly sounds peculiar. We will analyze Robinson’s passage in its context in Section 1.4.
1.2. Gaifman’s reading of Robinson
Erhardt thanks Gaifman for encouraging him to read Robinson’s text Formalism 64 [Erhardt 2024, p. 17]. Gaifman may have been the original source of a characterisation of Robinson as a finitist. Indeed, Gaifman claims in 2012:
Abraham Robinson, who was a finitist, or something very near to it, realized the seriousness of the limitations that his position implied with regard to syntactic concepts that required quantification over infinite domains. [Gaifman 2012, p. 488; emphasis added]
Gaifman’s own position is analyzed in Section 4.5. Was Robinson a finitist as claimed by Gaifman, and does Robinson’s position entail serious ‘limitations’ as Gaifman claims? To address the issue, it is crucial to distinguish between finitism in a narrow sense and finitism in a broad sense. In its broad sense, finitism denotes opposition to the use of infinitary concepts at the metamathematical level (as in Hilbert’s program). In its narrow sense, finitism is characterized by an opposition to the use of infinite totalities at both the metamathematical and the mathematical level. Thus, many intuitionists and constructivists were opposed to the use of certain infinite totalities in mathematical practice.
Erhardt’s claim of tension in Robinson’s work stems from a conflation of these two meanings of finitism. The fact that Robinson’s position is more accurately described as Formalism than finitism is evident from his recommendation concerning the business of mathematics:
[W]e should continue the business of Mathematics “as usual,” i.e., we should act as if infinite totalities really existed.
[Robinson 1965, p. 230; emphasis added]
Finitists (in the narrow sense) never “act as if infinite totalities really existed.” The idea that Robinson was not a finitist in the narrow sense is not merely a matter of our opinion against Erhardt’s. Indeed, Robinson explicitly argued against some finitists’ rejection of the use of infinitary terms in mathematics:
Those who adopt this attitude [including the Intuitionists] think that a concept, or a sentence, or an entire theory, is acceptable only if it can be understood properly and that a concept, or sentence, or a theory, is understood properly only if all terms which occur in it can be interpreted directly, as explained. By contrast, the formalist holds that direct interpretability is not a necessary condition for the acceptability of a mathematical theory. [Robinson 1965, p. 234; emphasis in the original]
One such intuitionist is Dummett, who claimed the following:
Constructivist philosophies of mathematics insist that the meanings of all terms, including logical constants, appearing in mathematical statements must be given in relation to constructions which we are capable of effecting, and of our capacity to recognise such constructions as providing proofs of those statements; … [Dummett 1975, p. 301; emphasis added]
Robinson clearly disagreed with Dummett’s claim that all terms must be assigned such a direct meaning. Robinson concluded:
To sum up, the direct interpretability of the terms of a mathematical theory is not a necessary condition for its acceptability; a theory which includes infinitary terms is not thereby less acceptable or less rational than a theory which avoids them. [Robinson 1965, p. 235; emphasis added]
Here Robinson endorsed the Formalist position and moreover contrasted it with finitism in the narrow sense.
1.3. Standard model
The source of Platonists’ discomfort with Formalism (and their proclivity to paint Formalists as finitists) is identified in section 10 of Robinson’s text Formalism 64:
In particular, I will mention here the assumption that there exists a standard or intended model of Arithmetic or (alternatively, but relatedly) of Set Theory. Clearly, to the formalist, the entire notion of standardness must be meaningless, in accordance with our first basic principle.333[Robinson 1965, p. 242]; emphasis in the original. For a discussion of this use of the term standardness see Section 4.1.
As discussed in Section 1.1, Robinson denied the existence of a referent for such a standard model (a.k.a. intended interpretation) of $\mathbb{N}$ or $\mathbb{R}$, in either the physical or any Platonic realm. To Platonists, this may appear as a narrow finitist stance, but they may well ponder why Robinson did not name his article “Finitism 64”.
Erhardt goes on to comment as follows:
There is, however, another alternative: more strongly committing to finitism. This is the path taken by various constructivists, who rather than play the uninterpretable game of symbols choose, in a philosophically principled manner, to adopt a weaker form of mathematics. [Erhardt 2024, p. 15; emphasis added]
Here Erhardt may be right about constructivists (rather than Robinson) describing classical mathematics as an ‘uninterpretable game of symbols’, [^4] Thus, Bishop writes: “The successful formalization of mathematics helped keep mathematics on a wrong course. … Mathematics becomes the game of sets, which is a fine game as far as it goes, with rules that are admirably precise” [Bishop 1967, p. 4]. In the same vein: “If every mathematician occasionally, perhaps only for an instant, feels an urge to move closer to reality, it is not because he believes mathematics is lacking in meaning. He does not believe that mathematics consists in drawing brilliant conclusions from arbitrary axioms, of juggling concepts devoid of pragmatic content, of playing a meaningless game” [Bishop 1967, p. viii]. though Erhardt does not admit in his article that Robinson himself does not describe it in these terms. Erhardt’s first sentence is based on the unfounded assumption that Robinson was a finitist, as we have already discussed; Erhardt compounds his error - of attributing to Robinson the position of narrow finitism [^5] As is evident from Erhardt’s comment on Robinson’s would-be reaction to Wiles’ proof of Fermat’s Last Theorem; see Section 1.5. - by taking it upon himself to offer advice as to how to practice the latter.
1.4. Robinson’s passage in context
In Section 1.1, we saw that Erhardt claimed that Robinson was a finitist based on a passage from Robinson’s text Formalism 64. Examining Robinson’s passage in context reveals a rather different picture. Robinson begins by mentioning the traditional positions of mathematical philosophy (Formalism, Intuitionism, Logicism) on page 228 of his text Formalism 64. On page 230, he mentions a philosophical school of nominalism. [^6] Today the term nominalism may refer to any anti-Platonist philosophy of mathematics. Robinson uses the term in a narrower sense. Robinson does not treat nominalism in much detail, on the grounds that to him it represents “little depth from the mathematical point of view” [Robinson 1965, p. 231]. Robinson claims that to a nominalist, there is not much difference between a set of five objects and an infinite set. Both are ‘illusory’ to a nominalist, according to Robinson:
To a nominalist, the existence of a set of five elements is no less illusory than the existence of the totality of all natural numbers. At the other end of the scale are the so-called platonic realists or platonists who believe in the ideal existence of mathematical entities in general, including the existence of transfinite sets of arbitrarily large cardinal numbers to the extent to which they can be introduced at all by means of suitable axioms. [Robinson 1965, p. 230; emphasis added]
By the time Robinson gets to the “five particular chairs” (as quoted by Erhardt) at the bottom of page 230, it is clear that his goal is to stress the difference between his position and that of nominalists. After noting the difference between “five particular chairs” and an infinite set, Robinson emphasizes on page 231 that describing infinite totalities as ‘meaningless’ does not mean that “such a theory is therefore pointless or devoid of significance” [Robinson 1965, p. 231].
As noted in Section 1.1, Robinson clarified his adjective ‘meaningless’ in his 1975 text where he stated that what he has in mind is an absence of a reference; namely, the term infinite totality does not refer to anything in either the real or a Platonic realm. Already in Formalism 64 he emphasized that he rejected the idea that infinite totalities exist “either really or ideally” [Robinson 1965, p. 230]. Here ‘really’ refers to the physical realm, and ‘ideally’ to a Platonic realm, as he states explicitly in the passage quoted above. It emerges that Erhardt’s presentation of Robinson’s position, as when Erhardt claims that it “places Robinson at odds with typical mathematical practice” [Erhardt 2024, p. 3], conflates the generic meaning of terms such as to grasp and meaningless, with the precise technical meaning attributed to such terms by Robinson.
1.5. Fermat’s Last Theorem: is the proof legitimate?
The conflation of the generic and the technical meaning of the term meaningless has a further effect of leading Erhardt to a preposterous misrepresentation of Robinson’s position, as when Erhardt claims:
Though Fermat’s Last Theorem is a general result proven by illegitimate methods and that purports to say something about all natural numbers—constituting an illicit reference to actual infinity—one can derive from it a potentially infinite number of legitimate, material claims. Yet the fact that the proof is illegitimate on Robinson’s view is crucial to our appraisal of his view. [Erhardt 2024, p. 12; emphasis added]
Erhardt’s illegitimacy claim stems from his mistaken identification of Robinson as a narrow finitist. But Wiles’ proof of Fermat’s Last Theorem would not be ‘illegitimate on Robinson’s view’ as Erhardt claims. While Wiles’ proof may seem illegitimate to those finitists who view proofs involving infinite totalities as meaningless in the generic sense of the term, it is certainly legitimate on Robinson’s view, collapsing what Erhardt describes as “our appraisal of his view” due to Erhardt’s conflation of distinct meanings of the term meaningless.
While discussing Robinson’s commitment to potential infinity, Erhardt claims that Robinson is a ‘devout realist’ about each natural number:
In opposition to nominalists (and a fortiori, fictionalists), Robinson is a devout realist with regard to each individual natural number. This is evidenced by his commitment to potential infinity, …777[Erhardt 2024, note 22]; emphasis on “devout realist” added.
Robinson himself would have likely rejected such a description of his position. Being able to grasp a concept (such as a potential infinity of metalanguage integers) does not commit one to a reality of a Platonic object. It emerges that Erhardt is trying to foist an untenable position on Robinson.
With regard to the issue of nominalism and fictionalism mentioned by Erhardt above, we note a pecularity of Erhardt’s approach. Erhardt is keen to cast Robinson as an instrumentalist, rather than one who regards infinite classes as useful fictions. Erhardt dismisses the suggestion that Robinson was a fictionalist, in spite of acknowledging that Leibniz, an avowed fictionalist, “had a profound effect” on Robinson [Erhardt 2024, p. 8 and note 22]. The rationale for such a dismissal is most peculiar: Erhardt considers that it would be anachronistic for Robinson, in 1964, to see himself as a fictionalist since Field’s ‘introduction of fictionalism’ occurred only in 1980 (see [Field 1980]). Erhardt is right to note that Field’s fictionalism, a form of nominalism, conflicts with Robinson’s repeated rejection of nominalism. But such a conflict arises only if fictionalism is assumed to be necessarily a nominalist proposal. Leibniz’s fictionalism offers an alternative to nominalism. Leibniz treated infinitesimals, along with negative and imaginary numbers, as well-founded fictions; see [Sherry and Katz 2012]. Although he sometimes suggested that infinitesimals are eliminable in the manner of Archimedean exhaustion arguments, he made no such suggestion for negatives and imaginaries. In all three cases, it is the contribution to systematicity that establishes the mantle ‘well-founded fiction’, rather than any sort of nominalist reduction.
1.6. Symptomatic keyword list and imaginary tensions
Already Erhardt’s keyword list is symptomatic of a problem with his text: the keyword finitism appears, but formalism does not. Erhardt misinterprets Robinson’s comment about being ‘unable to grasp’ actual infinite collections, by viewing it as a finitist stance. As analyzed in Section 1.4, Robinson merely sought to distance himself from nominalism and Platonism. Erhardt’s Abstract claims to detect a ‘tension’ between Robinson’s philosophy and his mathematical practice:
The foundational position he inherited from David Hilbert undermines not only the use of nonstandard analysis, but also Robinson’s considerable corpus of pre-logic contributions [^8] Erhardt’s wording here constitutes a historical error. He makes it appear as though Robinson’s work in aeronautics, such as his 1956 book Wing Theory [Robinson and Laurmann 1956], preceded his work in logic. This is incorrect. For example, Robinson’s dissertation The metamathematics of algebraic systems [Robinson 1949], advised by Paul Dienes, dates from 1949; see https://www.genealogy.math.ndsu.nodak.edu/id.php?id=15886. See also Dauben [Dauben 1995, p. 157]. to the field in such diverse areas as differential equations and aeronautics. This tension emerges from Robinson’s disbelief in the existence of infinite totalities … [Erhardt 2024, Astract; emphasis added]
Erhardt is referring to the book Non-standard Analysis [Robinson 1966]. However, Erhardt’s argument for his claim is too strong because it would apply to all Formalists. Erhardt’s Abstract accuses Robinson of “giving up on a commitment to reconciling” his philosophy and his mathematical practice. Possibly a Platonist would tend to view all Formalists in such a fashion. This may betray Erhardt’s own philosophical stance.
In sum, Erhardt’s depiction of Robinson’s philosophical position falls prey to a straw man fallacy. Erhardt’s claim in his abstract that Robinson’s philosophical position is somehow at odds with his work in applied mathematics, is unfounded; see further in Section 4.1.
2. Twin prime conjecture and types of infinite totalities
In his introduction, Erhardt claims that
[E]ven basic conjectures of number theory, such as the twin prime conjecture, presuppose the existence of infinite totalities. [Erhardt 2024, p. 2]
However, such a claim involves a conflation of distinct meanings associated with the term infinite totality; namely, of different levels of language or theory. The twin prime conjecture (TPC) can be formulated in Peano Arithmetic (PA). This can be done, for example, by the formula on page 13, line 4 in [Erhardt 2024], which we will express as follows:
$\forall x\,\exists y\,\phi(x,y),$ (2.1)
where the formula $\phi$ says that $y>x$ and both $y$ and $y+2$ are prime. Note that PA is a theory that does not know any infinite sets and that is even biinterpretable with the theory obtained from ZF by replacing the axiom of infinity by its negation, and adding $\in$-induction; see [Ackermann 1937]; [Kaye and Wong 2007]. The claim that the TPC presupposes the existence of infinite totalities is incorrect with respect to the object theory PA in which the TPC is formulated (because PA does not know any infinite totalities).
On the other hand, such a claim is true with respect to a metatheory in which PA is interpreted in an appropriate structure and thus given a semantics. On this reading, the term infinite totality refers to quantification over an infinite domain. Since formulation (2.1) involves such quantification, it can be said to involve infinite totalities in such a different sense. The special status of $\Sigma^{0}_{1}$ and $\Pi^{0}_{1}$-sentences is discussed in Section 4.5.
With regard to the first sense of the term infinite totality, note that it is not the twin prime conjecture that presupposes the existence of infinite totalities, but rather the assumption that there must be a determinate answer to whether it is true or false, based on the belief that the entire universe of all natural numbers exists as a completed infinity somewhere. Formalists, including Robinson, refrain from making such assumptions.
3. Fallacies and misrepresentations
We document several fallacies and misrepresentations in Erhardt’s text.
3.1. A logical fallacy
In Formalism 64, Robinson presents his disagreement with his Platonist opponents in the form of an Alice/Bob-type exchange [Robinson 1965, p. 231–232]. Erhardt misrepresents Robinson’s Alice/Bob-type presentation by attributing Alice’s position to Robinson himself, and failing to mention the more substantive of the two Alice/Bob issues. To elaborate, on page 231, Robinson mentions two objections that Alice (the opponent) might raise:
  1. (i)
    the ‘superior brain’ argument (which Robinson quickly dismisses), and
  2. (ii)
    the ‘postulation of physical/Platonist infinity’ argument.
Alice’s ‘superior brain’ argument (i) involves the idea that, while Robinson’s brain may have limitations ruling out a “clear conception of all sorts of infinite totalities,” his opponent may possess a superior conception of the said totalities.
Objection (ii) involves a postulation of infinite totalities in either the physical or a platonic realm. [^9] For the sake of completeness, we reproduce the full passage from Robinson: “An opponent to my position might put forward the following arguments. (i) He might say that I am unable to grasp the idea of an actual infinite totality merely because my brain suffers from a peculiar limitation. He might argue that he, on the contrary, has a clear conception of all sorts of infinite totalities or, at the very least, of the totality of natural numbers. (ii) Alternatively, my opponent may concede that he, also, is unable to grasp the idea of an infinite totality. But he may say that this does not in any way prove that infinite totalities do not exist. In order to show that they do exist he may appeal to the physical world. Or, if he does not wish to, or feels that he cannot, appeal to the physical world, he may affirm the existence of a platonic world which contains infinities of all sorts, or of some sort.” [Robinson 1965, p. 231].
Robinson quickly dismisses the first objection, pointing out that the postulation of such superiority “does not permit any further direct debate of the issue,” (Ibid.) and proceeds to comment on the second objection. Neither the noun psychology nor the adjective psychological occurs in Robinson’s discussion of either objection, or for that matter elsewhere in his text Formalism 64.
Surprisingly, Erhardt attributes Alice’s argument (i) to Robinson himself, and misleadingly presents the “no further direct debate” comment as Robinson’s final word in the issue. Writes Erhardt:
While Robinson has no difficulty grasping the possibility of arbitrarily large finite sets, which motivates his acceptance of the natural numbers taken individually, he reports that it is simply a psychological fact that he cannot grasp infinitary objects, obviating the possibility of finding common ground with disputants.101010[Erhardt 2024, p. 2]; emphasis on ‘grasping’ in the original; emphasis on ‘reports’ and ‘psychological’ added.
But Robinson ‘reports’ no such thing. As we already mentioned, the ‘psychological’ thing is a fabrication.
3.2. Robinson’s letter to Gödel
Erhardt’s discussion of Robinson’s letter to Gödel contains further misrepresentations:
In addressing a claim that someone with Gödel’s outlook might make–that he can fathom the infinite even if Robinson himself cannot–Robinson states that this objection ‘does not permit any further direct debate of the issue’. (Ibid.)
Contrary to what Erhardt’s quotation (taken out of context) suggests, Robinson is perfectly willing to engage in a debate - of the pertinent objection (ii) (see Section 3.1).
In his footnote 4, Erhardt similarly misrepresents Robinson’s comment “Hier stehe ich, ich kann nicht anders” in his letter to Gödel, [^11] This part of Robinson’s letter [Robinson 1973a] was not reproduced in [Gödel 2003]. by quoting it out of context. Robinson’s point is not to concede purported brain limitations but rather to reject Gödel’s realist assumptions. Moreover, Robinson gives his reasons for a reluctance to accept Gödel’s realist views concerning the hyperreals:
[T]he present evidence for the uniqueness of a non standard $\omega$-ultrapower of the reals is not strong.121212[Robinson 1973a]. Recall that hyperreal fields are often constructed as a quotient $\mathbb{R}^{\omega}\!/\mathcal{U}$ where $\mathcal{U}$ is a nonprincipal ultrafilter on $\omega$.
Erhardt’s attribution of Alice’s position to Robinson constitutes a logical fallacy [^13] Robinson is not the only mathematician whose position is misrepresented by Erhardt. Erhardt claims that Edward Nelson “believe[d] there is a largest number” [Erhardt 2024, p. 4]. To support his claim, Erhardt cites four texts by Nelson. We examined all four, including Nelson’s publication [Nelson 2011], but found no evidence of such a belief as claimed by Erhardt. Erhardt’s claim in [Erhardt 2024, note 7] that Nelson “believed Peano Arithmetic was inconsistent” is similarly inaccurate. Rather, Nelson suspected that PA was inconsistent. When he thought he had a proof he obviously thought it was inconsistent, but when Tao found an error in the proof, Nelson immediately accepted it. and does not inspire confidence in his analysis.
3.3. Verbal excesses from Kreisel to Erhardt
Similar remarks apply to certain verbal excesses, such as Erhardt’s claim that
Robinson has already admitted that such totalities are meaningless, and without an argument defending the use of these languages, formalism becomes an inconsistent project [Erhardt 2024, p. 6; emphasis added].
Erhardt does not explain how exactly Formalism would become an ‘inconsistent’ project in the absence of an argument defending the use of infinite totalities. Where is the inconsistency? His claim appears to be untenable.
Kreisel published his “Observations on popular discussions of foundations” in 1971 [Kreisel 1971]. Reviewer James D. Halpern for Mathematical Reviews described Kreisel’s “Observations” as “an unrestrained attack on P. J. Cohen … and on A. Robinson” and noted that
Most readers interested in foundations will probably find the previously mentioned papers of Cohen and Robinson profitable reading, the disparagement of these in the paper under review notwithstanding. [Halpern 1971]
As noted by Robinson,
Bernays, in an article published in Dialectica, criticized my attitude in his usual gentle manner, while Kreisel had stated his disagreement with me previously, also in his usual manner. [Robinson 1973b, p. 515]
The reference is respectively to [Bernays 1971] and [Kreisel 1971]. In a critical response to Robinson’s text [Robinson 1969], Bernays claimed that “there is no fundamental obstacle to attributing objectivity sui generis to mathematical objects.” [Bernays 1971, p. 178; translation ours]. Robinson appears to have responded in 1975 by including such ‘objectivity’ of infinitary entities alongside ‘reference’ as claims that a Formalist would reject; see e.g., [Robinson 1975, p. 49].
In his note 3, Erhardt seeks to distance himself from Kreisel’s verbal excesses, and describes Kreisel’s attack as ‘bias’ and ‘prejudice’. It is therefore disappointing to find Erhardt himself engaging in regrettable verbal excesses.
3.4. Whose game was it anyway?
Erhardt lodges the following claim concerning Robinson’s view of mathematics:
“He explains non-finitary mathematics as a collection of ‘uninterpretable games with symbols’ (Robinson 1969a, p. 47).” [Erhardt 2024, p. 10]
Looking up the original, one finds that Erhardt has applied a technique that is already familiar from Section 3.1: he attributes Robinson’s opponent’s position (in this case, ‘the intuitionist’) to Robinson himself. For the sake of completeness, we reproduce the full passage from Robinson’s text From a formalist’s point of view:
[T]he intuitionist may believe that the classical mathematican, whatever his underlying philosophy, is wasting his time in developing uninterpretable games with symbols, the sin of the formalist being the greater because he does so deliberately and consciously. [Robinson 1969, p. 47]
It emerges that, according to Robinson, it is the intuitionist (not Robinson himself) who is wont to make pejorative remarks about classical mathematicians and formalists allegedly developing ‘uninterpretable games with symbols.’ [^14] See note 4 for Bishop’s comments in this vein. Erhardt has again misrepresented Robinson’s position.
3.5. Platonism and consistency
Erhardt opens his text with the following sweeping claim concerning alleged realist attitudes among mathematicians:
[T]he mathematician tacitly adopts a realist outlook, imagining herself to investigate an independent realm of mathematical objects … about which she can discern objective truths. This assumption is crucial to the way mathematicians engage with their subject. [Erhardt 2024, pp. 1–2; emphasis added]
Erhardt’s ‘independent realm’ becomes an ‘independent reality’ by page 15. Postulating an ‘independent realm of mathematical objects’ is the gist of a Platonist position. Thus Erhardt postulates that being a Platonist is ‘crucial’ for the working mathematician. Or is it? Sir Michael Atiyah, the recipient of both a Fields Medal and an Abel Prize who was evidently not unsuccessful in “engaging with his subject”, had the following to say concerning independent realms:
The idea that there is a pure world of mathematical objects (and perhaps other ideal objects) totally divorced from our experience, which somehow exists by itself is obviously inherent nonsense. [Atiyah 2006, p. 38].
Erhardt’s opening remarks set the tone (and level of seriousness) for his 20-page text. In the context of a discussion of independence results (such as the Continuum Hypothesis), Erhardt claims that
[Robinson] does not consider the possibility that even if there are objective answers, they may be beyond our ability to know. [Erhardt 2024, p. 3; emphasis added]
But Robinson does in fact consider such a possibility, in the following terms:
If $X$ is the continuum hypothesis, …, then we know from the complementary results of K. Gödel and P. Cohen that both $X$ and non-$X$ are compatible with all known “natural” assumptions regarding the universe of sets (to use platonic language). While this suggests to the formalist that the entire notion of the universe of sets is meaningless (in the sense indicated by our first principle) the platonist merely concludes that the basic and commonly accepted properties of the universe of sets which are known to us at present are insufficient to decide the continuum hypothesis one way or the other. He will maintain, in this and similar cases, that at any rate only one of the alternatives that offer themselves is the correct one, i.e., is in agreement with the truth. [Robinson 1965, p. 232; emphasis added]
Robinson then goes on to analyze the difficulties of the Platonist position with regard to independence results. Thus, Erhardt’s claim is factually incorrect.
On page 15, Erhardt repeats the unfounded accusation aimed at Robinson, of “deflating mathematics to a game” [Erhardt 2024, p. 15]. There follows a curious paragraph on page 15 in Erhardt concerning Platonism and the problem of consistency:
All of this constitutes Robinson giving up on a serious commitment to his finitism. This is not to say that he should have pursued a definitive argument against platonism, but that there are more and less philosophically sophisticated ways of reconciling mathematical practice with the assertion that we cannot grasp infinite collections. (Ibid.; emphasis added)
Does Erhardt believe that Platonism provides a way out for the problem of consistency? The paragraph continues:
As it stands, suggesting that mainstream mathematics continue is tantamount to an admission, despite a reliance on infinitary notions, that Robinson believes it is consistent.
This is an allusion to Gödel’s second incompleteness result, asserting the impossibility of proving Con(PA) (as formalized by Gödel) within Peano Arithmetic itself (and similar results for stronger systems):
If not–for instance, if nonstandard analysis led to contradiction–then anything could be proven from it.
But such a conclusion would surely hold for just about any piece of mathematics, and of any philosophical stance. Or would it? Erhardt continues:
Robinson does not venture to explain why he takes infinitary mathematics to be consistent, though a compelling answer might have provided a rationale for accepting the finitary results it produces. In the absence of such an explanation, relying upon infinitary mathematics is itself a source of great risk, for there is no assurance that the derived concrete implications are valid. [Erhardt 2024, p. 15; emphasis added]
Erhardt holds that relying on infinitary mathematics is a source of great risk. The paragraph suggests that Erhardt believes that Platonism is somehow capable of eluding such a ‘risk’. The reasoning presumably is that if a Platonic realm of mathematics exists, of necessity it could be neither contradictory nor undecidable, as Erhardt goes on to comment on issues of faith:
Robinson is here relying on faith … if not about the universe of sets, then about basic facts concerning the natural numbers. This hunch is not a sufficient warrant to stand in place of a satisfying mathematical result, like the completion of Hilbert’s program, and so his instrumentalism is unjustified. (Ibid.; emphasis added)
We leave it to the reader to judge which position involves a greater leap of faith: that of relying on infinitary mathematics without a commitment that it possesses a reference (in the sense explained in Section 1.1), or that of placing one’s faith in a Platonic heaven to ensure the consistency of infinitary mathematics.
Exactly halfway through his text (on page 10 out of 20), Erhardt finally admits that he is familiar with the philosophical concept of not referring, when he describes Robinson’s approach to non-finitary mathematics as
[An] activity that demands no ontological commitments because so many of the terms in its expressions do not refer.
[Erhardt 2024, p. 10].
Note that Erhardt’s clause “do not refer” occurs nowhere in Robinson’s texts, and therefore constitutes Erhardt’s own paraphrase of Robinson’s position. This particular claim of Erhardt’s happens to be an accurate description of Robinson’s position concerning the lack of reference (in the sense detailed in Section 1.1), [^15] To be sure, symbols such as $\mathbb{N}$ and $\mathbb{R}$ can refer to the corresponding mathematical entities in the context of a traditional set theory such as ZF. However, Robinson specifically speaks of absence of reference to entities in the physical or a putative Platonic realm. unlike Erhardt’s claims regarding Robinson’s purported finitism.
3.6. Grand justification wherefore?
Erhardt finds fault with Robinson’s strategy as follows:
[Robinson] neglects to pursue a grand justification for infinitary mathematics by–in lieu of attempting to offer ‘a satisfactory intellectual motivation’–deflating mathematics to a game. [Erhardt 2024, p. 15; emphasis added].
The fact that Erhardt’s game claim involves a misattribution has already been addressed: Robinson didn’t say it; rather, he mentioned the intuitionist’s use of this pejorative term to criticize both classical mathematicians and formalists (see Section 3.4). Let us now examine Erhardt’s claimed ‘neglect’ of ‘grand justification’. The fact is not merely that Robinson had no such intention, but that moreover such a project would have been contrary to his philosophical stance. Erhardt fails to appreciate the pragmatic nature of Robinson’s philosophical stance, as we now elaborate.
In his 1969 text [Robinson 1969], Robinson amplifies his list of two items that he already presented in his Formalism 64: (i) ‘infinite totalities’ lack reference (in the sense explained in Section 1.1), and (ii) useful fictions: a mathematician should continue acting as if they really existed.
In 1969 Robinson adds a third item: (iii) against bottomless pragmatism: there is an identifiable core of logical thought that is common to all of mathematics. Even more importantly, Robinson uses these three items to eliminate what he diagnoses as philosophical dead-ends: items (i), (ii), and (iii) eliminate the positions of, respectively, the platonist, the intuitionist, and the logical positivist.
It emerges that Robinson is decidedly not in the business of developing ‘grand justification’ schemes. He is merely ruling out what seem to him to be common philosophical misconceptions, to avoid ending up in a philosophical dead-end. To reproach Robinson for not pursuing a grand justification scheme is to miss entirely the spirit of Robinson’s philosophical posture.
Erhardt claims that
While Robinson repudiates actual infinity, the mathematics of his career revolves around it. This incongruity mirrors the move from (i) to (ii); … Robinson says the same, noting that his position suffers a serious drawback because ‘the gap due to the absence of consistency proofs for the major mathematical theories appears to be inevitable and we have learned to live with it’. [Erhardt 2024, p. 10; emphasis added]
But there is no ‘incongruity’ here, and Robinson certainly did not “say the same” as Erhardt claims. This is yet another misrepresentation of Robinson’s position, involving a quotation out of context. The full quotation makes it clear that Robinson viewed ‘the absence of consistency proofs’ as a difficulty common to all schools in the Philosophy of Mathematics, and not merely ‘his position’ as Erhardt claims. Thus, Robinson wrote:
… all known positions in the Philosophy of Mathematics, including my own, still involve serious gaps and difficulties. Among these, the gap due to the absence of consistency proofs for the major mathematical theories appears to be inevitable and we have learned to live with it. [Robinson 1965, p. 236; emphasis added]
The inevitability of such a gap has been challenged in [Artemov 2024]. Some logicians are skeptical about the philosophical implications of Artemov’s result. Regardless of the outcome of that particular debate, there seems to be little reason to assume that incompleteness is less of a problem for Platonism than for Formalism, and Robinson certainly did not claim such a thing.
4. Types of standardness, realism, knowledge and truth
We examine some issues related to types of standardness, as well as some realists’ take on the relation of knowledge and truth. We take a closer look at some conflations of terms that lead Erhardt to dubious conclusions about Robinson’s work. We also examine whether the realism proposed by Erhardt and Gaifman has more convincing answers than Formalism when it comes to questions of mathematical knowledge and applicability.
4.1. Meanings of Standardness
As already mentioned at the end of Section 1.4, Erhardt appears to have difficulty keeping apart the generic and the technical meaning of terms like meaningless and to grasp. Following a discussion of Skolem’s nonstandard models and Robinson’s nonstandard analysis, involving the technical distinction between standard and nonstandard numbers, Erhardt writes:
“Adding to the irony–and irresolvable tension in Robinson’s work and philosophy–is that, due to its predication on the transfinite, Robinson himself says that ‘the entire notion of standardness must be meaningless […]’ (Robinson 1964/1979b, p. 242).” [Erhardt 2024, note 33; emphasis added]
The irony, according to Erhardt, is that Robinson himself talks ‘standard and nonstandard’, and then comes out swinging against ‘the entire notion of standardness’. However, the context - not clarified by Erhardt - of Robinson’s comment on ‘the entire notion of standardness’ is a discussion of the Platonist idea of a standard or intended model of arithmetic or set theory, as analyzed in Section 1.3. To reject such an idea, Robinson used the term ‘standard’ in its generic sense, whereas Erhardt misleadingly presents it as if what is involved is the technical sense used in nonstandard analysis.
In more detail, theories of nonstandard analysis exploit a mathematical concept called the standardness predicate, which distinguishes between standard and nonstandard numbers (and more general sets). Since this distinction can be thought of as a formalisation of Leibniz’s distinction between assignable and inassignable numbers, [^16] See e.g., the publications [Katz and Sherry 2013]; [Katz and Kuhlemann 2023/5]; [Katz et al. 2024]; [Ugaglia and Katz 2024], [Kuhlemann 2025]. the predicate could be referred to also as the assignability predicate. Meanwhile, Platonists believe in (and Robinson rejects) the existence of standard models a.k.a. intended interpretations of ${\mathbb{N}}$, ${\mathbb{R}}$, and other mathematical entities. When the issue is translated into the terminology of the assignability predicate and the intended interpretation, it is obvious that there is no connection between them, contrary to the impression Erhardt seeks to create.
4.2. Resolving tensions
It is worth examining Erhardt’s claim concerning an alleged ‘irresolvable tension in Robinson’s work and philosophy’ in more detail. Here one needs to distinguish between doing mathematics and talking about mathematics.
(1) The question whether $10000000000000001$ is a prime concerns a concrete natural number. One does not need Peano Arithmetic, let alone the infinite set ${\mathbb{N}}$, for it to be meaningful.
(2) On the other hand, to make sense of the question whether there are arbitrarily large numbers $n$ for which $110^{n}+$ 1 is prime does require some theory of arithmetic, and perhaps (if Erhardt’s reading of Quine is right) also an infinite set ${\mathbb{N}}$.
(3) The notions of formal logic such as variable, term, formula and proof are defined by recursive rules similar to the definition of natural numbers. We therefore have an analogous distinction to that between items (1) and (2) above, as follows.
(4) When we write down (more/less formal) actual proofs in some area of mathematics, these proofs are concrete objects analogous to concrete numbers such as $10000000000000001$. We do not need the collection of all variables, terms, formulas or proofs to verify that a given concrete proof really is a correct proof. This is what formalists like Robinson are doing when proving results in some area of mathematics, including aeronautics, they are working in; no actual infinity is necessarily involved here.
(5) Formalists can also be interested e.g., in determining what is provable. Then they are doing metamathematics, i.e., the area they are working in is logic. For this purpose one will likely want to exploit the kind of infinite collections listed in item (3).
Thus, contrary to Erhardt’s claim, there is no tension between Robinson’s work in aeronautics (mathematics), on the one hand, and his work in - or philosophy of - metamathematics (logic), on the other. One could make the following additional points.
(6) When mathematicians practice some area of mathematics, they prove theorems. Such theorems (for example, Wiles’s proof of Fermat’s Last Theorem) are concrete objects. No theory of proofs, and in particular no commitment to infinite collections, is needed to generate such proofs.
(7) A formalist mathematician may feel compelled (but, arguably, is not required qua philosopher) to study mathematical practice mathematically. If and when doing that, they practice mathematics in another field: logic, proof theory, model theory… In such practice, they may find (referenceless) infinite collections of variables, formulas or proofs useful. The remark of item (6) applies to such practice, as well. There is no need for actual infinity in order to give a proof of and find useful Gödel’s Incompleteness theorem. Thus even in logic there are results that do not depend on a commitment to any actual infinity.
In sum, there is neither contradiction nor tension between a formalist’s practice and theory (philosophy).
4.3. Varieties of realism
Not all mathematical realists hold identical views. We will examine Erhardt’s position in a framework proposed by Hamkins. In an influential 2012 article, Hamkins provides the following useful perspective:
The multiverse view in set theory … is the view that there are many distinct concepts of set, each instantiated in a corresponding set-theoretic universe. The universe view, in contrast, asserts that there is an absolute background set concept, with a corresponding absolute set-theoretic universe in which every set-theoretic question has a definite answer. The multiverse position, I argue, explains our experience with the enormous range of set-theoretic possibilities, a phenomenon that challenges the universe view. [Hamkins 2012, Abstract]; see also [Hamkins 2021, Section 8.14]
Hamkins goes on to state his position with regard to the Continuum Hypothesis (CH):
In particular, I argue that the continuum hypothesis is settled on the multiverse view by our extensive knowledge about how it behaves in the multiverse, and as a result it can no longer be settled in the manner formerly hoped for. (Ibid.)
In more detail, Hamkins views CH as a “switch” that can be turned on and off as it were, by passing to a suitable forcing extension of the current instance of a set-theoretic universe.
Hamkins considers himself a realist of the multiverse, thereby discarding the assumption of the uniqueness of a set-theoretic universe as sometimes held by other realists.
We will now examine Erhardt’s stance, in the context of Hamkins’ dichotomy. Erhardt writes:
[T]he mathematician tacitly adopts a realist outlook, imagining herself to investigate an independent realm of mathematical objects–one that includes infinite totalities like the natural numbers–about which she can discern objective truths.
[Erhardt 2024, p. 1].
While this passage suggests that Erhardt’s sympathies are with the realist, it does not indicate whether or not this constitutes realism about a unique set-theoretic universe. Reading further, we find the following comment concerning Zermelo–Fraenkel set theory with the Axiom of Choice (ZFC) and the CH:
Consider Cantor’s continuum problem: Is there an intermediate cardinality between the naturals and the reals? If you are a realist about infinite totalities, then the continuum problem is an obvious question to ask. However, ZFC+CH and ZFC+CH [sic] [^17] The second occurrence of ZFC+CH is a typo and should be “ZFC+$\neg$ CH”. are both consistent, which Robinson takes as implying that there must be no universe of sets. A platonist will respond that we have simply not yet found the axioms stating the basic properties of the universe of sets sufficient to decide CH one way or the other. This is a valid objection, but it does not sway Robinson. [Erhardt 2024, note 5; emphasis added]
Here Erhardt is sympathetic to the idea that CH ought to possess a definite truth value yet to be determined, and considers such reasoning as ‘valid’. Erhardt’s position concerning CH is clearly at odds with Hamkins’ multiverse-realism, and specifically with Hamkins’ position with regard to CH (namely, that it need not possess a definite truth value yet to be determined). Our tentative conclusion is that Erhardt is a universe-realist. [^18] Kreisel asks: “Can we extend the language of set theory to … decide CH by means of axioms (which are evident for the intended interpretation) of the extended language?” [Kreisel 1971, pp. 196–197]. Thus far, evidence in favor of an affirmative answer has been weak. As argued above, this may color his evaluation of Robinson’s philosophical position.
4.4. Applicability of mathematics
The far-reaching applicability of mathematics in diverse areas is certainly a famous problem without easy answers. However, when Erhardt claims that
[Robinson’s] foundational position undermines not only the use of nonstandard analysis, but also Robinson’s considerable corpus of pre-logic contributions [^19] See note 8. to the field in such diverse areas as differential equations and aeronautics [Erhardt 2024, Abstract]
one may well wonder how exactly the postulation of a Platonic set-theoretic universe would help explain the applicability of mathematics in “such diverse areas as differential equations and aeronautics.” In other words, how would the putative existence of a mind-independent realm of mathematical abstracta help explain the ability of an embodied brain to apply the said abstracta to the solution of concrete problems of airplane wing design (cf. [Robinson and Laurmann 1956])? It would seem that the difficulty involved is no less challenging than the applicability of what Leibniz already referred to as useful fictions. [^20] See [Katz and Sherry 2013]; [Katz et al. 2024]; [Ugaglia and Katz 2024]; [Kuhlemann 2025].
Robinson proposed an apt diagnosis of the Platonist’s aversion to the Formalist’s position:
[I]t is perhaps natural that a mathematician should resent suggestions which deprive him of the comforting feeling that he, like the physicist or biologist, spends his life in the exploration of some form of reality. [Robinson 1969, p. 46].
As far as the criterion of the applicability of mathematics in the natural sciences, Robinson explains:
[T]his criterion does not imply that the terms of the theory should be interpretable directly and in detail. It is sufficient that we should have rules which tell how to apply certain relevant parts of our theory to the empirical world. [Robinson 1965, p. 234].
Thus, infinitary terms may be referenceless, but their applications can nonetheless be meaningful.
4.5. Gaifman’s realism
[Gaifman 2012] claims to use the term realism in the sense of realism in truth-value. In relation to the natural numbers, this means assuming that every PA sentence is a factual statement, i.e., has an objective truth value (even if it is undecidable in PA itself). As Robinson pointed out almost half a century earlier,
[A]s a matter of empirical fact the platonists believe in the objective truth of mathematical theorems because they believe in the objective existence of mathematical entities. [Robinson 1965, p. 230; emphasis in the original]
There is internal evidence in Gaifman’s article that such is indeed his position, as illustrated by his snide comment about Hamkins and Robinson:
Another possibility has emerged from the multiverse conception proposed by Hamkins (2011). On this view, the model of natural numbers depends on the set-theoretic universe containing the model. What it comes to is that we have a clear enough conception of models of PA, or of some extensions of PA, but we have no clear distinction between standard and non-standard models. Perhaps Abraham Robinson, the founder of non-standard mathematics, who was not a set theoretician, would agree to that. [Gaifman 2012, p. 489]
Gaifman’s professed truth-value realism appears to be consistent with the belief that the ‘standard or intended model of arithmetic’ is a meaningful notion. Recall that Robinson rejects such a notion (see Section 1.3).
Note that “$\phi$ is provable” and “$T$ is consistent” are respectively $\Sigma^{0}_{1}$ and $\Pi^{0}_{1}$ sentences (when encoded in PA). Gaifman apparently holds that one commits oneself to realism in truth value (in relation to the natural numbers) when one uses the terms provable and consistent in metamathematics. He seems to find it philosophically unacceptable to assume that metamathematical statements about the provability of a sentence or the consistency of a theory could be undecidable. Meanwhile, for a formalist practicing finitistic metamathematics in Primitive Recursive Arithmetic (PRA) such a posture toward undecidability is a natural consequence. The “finitistic semantics” of PRA is naturally incomplete; see [Tait 1981]. A $\Pi^{0}_{1}$ sentence is ‘true’ if it can be proved by induction; it is ‘false’ if one can provide a counterexample. A $\Sigma^{0}_{1}$ sentence is ‘true’ if one can provide an example; it is ‘false’ if one can provide an inductive proof of the corresponding $\Pi^{0}_{1}$ sentence expressing its negation; see further in [Simpson 2009].
Similarly to intuitionism, truth in this semantics is defined by provability, and falsity by refutability. Consequently, in metamathematics, on the basis of potential infinity, there is no tertium non datur. Such a position is unacceptable for Gaifman.
Such a metamathematics can nevertheless prove meaningful statements. For instance, as explained in [Hrbacek and Katz 2021] the conservativity of the theory SPOT over ZF can be proved in WKL0 (which is not finitistic). But this conservativity result is a $\Pi^{0}_{2}$ sentence, and according to the results of reverse mathematics [Simpson 2009], there is therefore a proof in PRA, where the result is formulated as $\phi(m,f(m))$ with a primitive recursive function. Such a formulation is finitistic.
The fact that realist positions (like Gaifman’s) make it harder to appreciate the hidden resources of the theory of the real number line was noted by Massas in the following terms:
[T]he conservativity results in [Hrbacek and Katz] [^21] Massas is referring to [Hrbacek and Katz 2021]. would likely fall short of convincing anyone who thinks that any existence claim regarding Robinsonian infinitesimals is simply false.
[Massas 2024, p. 263]
Paul Cohen seemed to regard it as a prerequisite for his progress in foundational research to abandon the idea that there is a unique (intended) interpretaion of $\mathcal{P}({\mathbb{N}})$ that conforms to our intuition:
I can assure that, in my own work, one of the most difficult parts of proving independence results was to overcome the psychological fear of thinking about the existence of various models of set theory as being natural objects in mathematics about which one could use natural mathematical intuition. [Cohen 2002, p. 1072]
4.6. Knowledge-transcendent truth?
With regard to the relation between independence results and mathematical realism, Gaifman mentions ‘knowledge-transcendent truth’:
If an independence result indicates that some mathematical truths outstrip our capacities of knowing them, then it points to knowledge-transcendent truth, hence to realism. [Gaifman 2012, p. 498]
Objectively speaking, the implication would seem to go the other way around: it is only if one wishes to defend realism that one may be led to postulate knowledge-transcendent truth. Such ‘knowledge-transcendent truth’ is mentioned twice in the article (the second occurrence is in the concluding section). It emerges that when Gaifman claims that something points toward realism, what is he has in mind is that it points toward transcendent truth. However, Gaifman’s argument appears to be circular. Namely, the starting assumption that “independence results indicate that some truths outstrip our capacity” itself depends on a realist standpoint. An anti-realist has little reason to assume that independence results would indicate such a thing. So the full Gaifmanian circularity would run as follows:
[…]
Gaifman claims further:
I am mostly concerned with independence results for first-order arithmetical statements. My general view is that these results do not affect our conception of the standard natural numbers. Hence, they reinforce any realism that is based on this conception. [Gaifman 2012, p. 509]
Gödel’s second incompleteness theorem as applied to PA is a classical arithmetic independence result. Gaifman appears to claim that such a result would “reinforce realism.” Accordingly, the independence of CON(PA) of PA would reinforce realism. One wonders what Gaifman would make of Artemov’s recent proof of CONS(PA) within PA.222222[Artemov 2024] proposed an alternative formalisation ConS(PA) of consistency following Hilbert. Giaquinto quotes Robinson in [Robinson 1965, p. 239] to the effect that “I cannot see at this time how a form of reasoning which attempts to escape the consequences of Gödel’s second theorem (such as Gentzen’s consistency proof or any other consistency proof for Arithmetic) can remain strictly finitistic and hence interpreted” and claims that “the formalist foundational programme is irredeemable” [Giaquinto 1983, p. 129]. Artemov’s result provides a possible way of escaping the consequences of Gödel’s second theorem, and therefore possibly undermines Giaquinto’s rash conclusion. Similar remarks apply to Giaquinto’s surprising claim that “the resuscitation of formalism by Robinson and Cohen (following Cohen’s proof of the independence of the Continuum Hypothesis) is untenable and barren” [Giaquinto 1983, p. 120]. If even independence reinforces realism, lack of independence (in the sense of Artemov’s result) would seem to also reinforce realism. This would suggest that Gaifman’s variety of realism is unfalsifiable.
All known undecidable first-order arithmetic sentences have a preferred truth value. For example, CON(PA) is generally considered to be true (and is of course provable in a stronger system such as ZFC), unlike higher-order statements such as CH. Some mathematicians view this as an argument in favor of realism about PA. However, such an argument does not appear in [Gaifman 2012].
5. Conclusion
While Erhardt goes on eventually to discuss Robinson’s mathematical Formalism, his opening presentation of Robinson’s position as finitism is off the mark, and his attempt to present Robinson’s opponent’s position as Robinson’s own does not inspire sufficient confidence to give credence to his speculations concerning Robinson’s analysis of potential infinity.
Erhardt’s endorsement of a naive realism about a standard model of ZFC, including his faith in a determinate truth value of CH (yet to be ‘discovered’), fails to take into account the recent literature in the philosophy of the foundations of mathematics, such as the multiverse perspective of [Hamkins 2012]; see also [Hamkins 2021, Section 8.14]. Due to his misreading of Robinson’s terms meaningless, to grasp, and standard, Erhardt sees an irony and an unresolvable tension between Robinson’s use of nonstandard models of the real numbers and his statement that “the entire notion of standardness must be meaningless.” However, there is neither irony nor tension here, since Robinson is referring to his previous statement that infinite sets, including both standard and non-standard models, lack reference. He did not mean that they are pointless or devoid of significance.
In his 1969 text, Robinson envisions
… a solution which involves the equal acceptance of several kinds of set theory. At this point, it is natural to recall the historical development of Euclidean geometry whose analogy with the recent development of axiomatic set theory is perhaps closer than is generally appreciated. [Robinson 1969, pp. 46-47; emphasis added]
The comparison between the existence of distinct set theories and the existence of distinct geometries was pursued in more detail by [Cohen and Hersh 1967]. Different set theories of this type were indeed developed shortly after Robinson’s passing. [Hrbacek 1978] and [Nelson 1977] independently developed axiomatic approaches to nonstandard analysis as an alternative to model-theoretic approaches. In fact, Robinson’s visionary insight is arguably more likely to emerge from a Formalist than a Platonist philosophical stance. [^23] Our conclusion here is at variance with Kreisel’s. Kreisel claimed the following: “having accepted the formalist scheme, one is not merely unwilling to appreciate (or study!) further work in foundations, but simply unable to do so without reexamining one’s ideas” [Kreisel 1971, p. 191] (emphasis in the original). However, Hrbacek and Nelson were indeed able to appreciate, study, and make progress in foundations without adopting Platonist ideas.
While traditional set theory is formulated in the $\in$-language, the axiomatic approaches enrich the language of set theory by means of the introduction of a one-place predicate ‘standard’ or ‘st’, and are thus formulated in the st-$\in$-language. Such theories are conservative over ZFC.  Suitable axioms govern the interaction of the new predicate with the traditional ZFC axioms. In the axiomatic approaches, infinitesimals are found within ${\mathbb{R}}$ itself (rather than in an extension, as in the model-theoretic approaches to nonstandard analysis), and nonstandard integers in ${\mathbb{N}}$ itself. This challenges Platonist notions about ${\mathbb{N}}$ and ${\mathbb{R}}$ as determinate (or even mind-independent) entities. Robinson’s Formalism remains a viable alternative to mathematical Platonism.
Acknowledgments
We are grateful to Karel Hrbacek for insightful comments.